# Federico Echenique and Farzad Pourbabaee

### v4

@@ 5.3 Concentration and Isoperimetric Inequalities
Assume towards a contradiction that $\bigcap_{i\in I}\mathcal{B}_{i}(f_{i})^{\delta}\neq\emptyset$, while $f$ is $\varepsilon$-Pareto dominated. Choose $\eta\in\bigcap_{i\in I}\mathcal{B}_{i}(f_{i})^{\delta}$. Since $f$ is $\varepsilon$-Pareto dominated, there is $g\in\mathcal{F}_{\bar{\omega}}$ such that $(1-\varepsilon)g_{i}\succ_{i}f_{i}$ for all $i$. By definition of the subjective belief set, and continuity one obtains that $\mu_{i}\cdot\big{[}(1-\varepsilon)g_{i}-f_{i}\big{]}>0$ for all $\mu_{i}\in\mathcal{B}_{i}(f_{i})$. Choose $\tilde{\mu}_{i}\in\operatorname*{arg\,min}\{\lVert\eta-\mu_{i}\rVert:\mu_{i}\in\mathcal{B}_{i}(f_{i})\}$. Observe that $\lVert\eta-\tilde{\mu}_{i}\rVert<\delta$, hence
@@
Since both $f$ and $g$ belong to $\mathcal{F}_{\bar{\omega}}$, the leftmost side above is equal to $-\varepsilon\bar{\omega}$, thus leading to a contradiction.∎
@@ 6.2 Proof of Proposition 3
Since $\Pi$ is a $(d-1)$-dimensional surface with constant width, then Schramm (1988) implies that
@@
The volume of the $(d-1)$-dimensional unit $\ell_{2}$ ball and the $(d-1)$-dimensional probability simplex $\Delta_{d}$ are respectively equal to $\mathsf{Vol}(\mathbb{B}_{2}^{d-1})=\pi^{(d-1)/2}/\Gamma\left(\frac{d+1}{2}\right)$ and $\mathsf{Vol}(\Delta_{d})=\sqrt{d}/\Gamma(d)$. Therefore,

### v5

@@ 5.3 Concentration and Isoperimetric Inequalities
Assume towards a contradiction that $\bigcap_{i\in I}\mathcal{B}_{i}(f_{i})^{\delta}\neq\emptyset$, while $f$ is $\varepsilon$-Pareto dominated. Choose $\eta\in\bigcap_{i\in I}\mathcal{B}_{i}(f_{i})^{\delta}$. Since $f$ is $\varepsilon$-Pareto dominated, there is $g\in\mathcal{F}_{\bar{\omega}}$ such that $(1-\varepsilon)g_{i}\succ_{i}f_{i}$ for all $i$. By definition of the subjective belief set, and continuity one obtains that $\mu_{i}\cdot\big[(1-\varepsilon)g_{i}-f_{i}\big]>0$ for all $\mu_{i}\in\mathcal{B}_{i}(f_{i})$. Choose $\tilde{\mu}_{i}\in\operatorname*{arg\,min}\{\lVert\eta-\mu_{i}\rVert\colon\mu_{i}\in\mathcal{B}_{i}(f_{i})\}$. Observe that $\lVert\eta-\tilde{\mu}_{i}\rVert<\delta$, hence
@@
Since both $f$ and $g$ belong to $\mathcal{F}_{\bar{\omega}}$, the leftmost side above is equal to $\varepsilon\bar{\omega}$, thus leading to a contradiction.∎
@@ 6.2 Proof of Proposition 3
Since $\Pi$ is a (d1)(d-1)-dimensional surface with constant width, then Schramm (1988) implies that
@@
The volume of the (d1)(d-1)-dimensional unit $\ell_{2}$ ball and the (d1)(d-1)-dimensional probability simplex $\Delta_{d}$ are respectively equal to $\mathsf{Vol}(\mathbb{B}_{2}^{d-1})=\pi^{(d-1)/2}/\Gamma\left(\frac{d+1}{2}\right)$ and $\mathsf{Vol}(\Delta_{d})=\sqrt{d}/\Gamma(d)$. Therefore,